OpenStudy (anonymous):

A recycling company recycles wet and dry wastes. In a week, each ton of dry waste takes 35 hours for collection and 55 hours for processing. Each ton of wet waste takes 42 hours for collection and 63 hours for processing. The company's collection workforce can work for maximum 150 hours in a week, and the processing machine can work for maximum 200 hours in a week. Let D represent the number of tons of dry waste that the company can recycle in a week and W represent the number of tons of wet waste that the company can recycle in a week.

OpenStudy (anonymous):

Which of the following statements best explains whether the company will be able to recycle 3 tons of dry waste and 2 tons of wet waste in a week? A. The company will be able to recycle these wastes in a week because of the system restrictions of 35D + 63W ≤ 200 and 55D + 42W ≤ 150. B. The company will be able to recycle these wastes in a week because of the system restrictions of 35D + 55D ≤ 150 and 42W + 63W ≤ 200. C.The company will not be able to recycle these wastes in a week because of the system restrictions of 35D + 42W ≤ 200 and 55D + 63W ≤ 150. D.The company will not be able to recycle these wastes in a week because of the system restrictions of 35D + 42W ≤ 150 and 55D + 63W ≤ 200.

OpenStudy (texaschic101):

dry waste (d), wet waste (w) 35d + 42w < = 150 55d + 63w <= 200 now....if dry waste(d) = 3 and wet waste (w) = 2 we sub 35(3) + 42(2) <= 150 105 + 84 <= 150 189 <= 150 (incorrect) 55(3) + 63(2) <= 200 165 + 126 <= 200 291 <= 200 (incorrect) answer D -- they won't be able to recycle the waste in a week because of the restrictions of 35d + 42w <= 150 and 55d + 63w <= 200.

OpenStudy (anonymous):

taco

OpenStudy (texaschic101):

burrito

OpenStudy (anonymous):

thank you @texaschic101